OpenStudy (anonymous):

At a carnival game, a large horizontal board has 20 holes of the same size cut in it. For the game, you toss a ball onto the board and it lands or rolls into one of the holes. The board is surrounded on all sides, so the ball is certain to fall in a hole. Three of the holes will earn a small prize and one of the holes will earn a large prize. If Grady buys a ticket for one ball that he'll toss on the board, what is the probability that he'll win a prize?\ 5% 15% 20% 40%

OpenStudy (anonymous):

help?

OpenStudy (angela210793):

I think 20

OpenStudy (anonymous):

i need 100%

OpenStudy (angela210793):

I thought...from 20 holes only in 4 of them u can win a prize.....so p=4/20*100=.2*100=20%

OpenStudy (anonymous):

thx!!

OpenStudy (angela210793):

I hope tht's the right answer